OpenStudy (anonymous):

(1/tan x)+tan x=sec x *csc x I have to prove one side of the equation is equal to the other. This is a proof.

OpenStudy (anonymous):

\[\frac{1}{\tan x}+\tan x=\cdots\\ \frac{\cos x}{\sin x}+\frac{\sin x}{\cos x}=\cdots\\ \frac{\cos^2x}{\sin x\cos x}+\frac{\sin^2x}{\sin x\cos x}=\cdots\\ \frac{\cos^2x+\sin^2x}{\sin x\cos x}=\cdots\] And so on

OpenStudy (anonymous):

thanks!